To break even in a manufacturing​ business, income or revenue R must equal the cost of production C. The cost C to produce x skateboards is c=63+22x. The skateboards are sold wholesale for $25 each, so revenue R is given by R=25x Find how many skateboards the manufacturer needs to produce and sell to break even.​ (Hint: Set the cost expression equal to the revenue expression and solve for​ x.)


1
Expert's answer
2022-01-17T15:57:25-0500

"\text{From the question, C$=63+22x$ and R$=25x$}\\\text{Then, since break even means (C)ost=(R)evenue, we have;}\\ \text{C$=$R}\Rightarrow63+22x=25x\Rightarrow63=3x\Rightarrow x=\frac{63}{3}=21.\\ \text{Hence, the manufacturer needs to produce and sell $\mathrm{21}$ \textrm{skate boards} to break even.}"